OpenStudy (anonymous):

Can someone tell me a formula to do these kinds of problems? Solve for x: 3x − 5 = 2x + 6. 1 −1 11 −11

OpenStudy (anonymous):

You want to start by getting all the x's on one side and all the terms w/o x on the other side

OpenStudy (decentnabeel):

\[3x-5=2x+6\quad :\quad x=11\]

OpenStudy (anonymous):

3x-2x= 6+5

OpenStudy (decentnabeel):

3x-5=2x+6 subtract 2x n both sides 3x-2x-5=2x-2x+6 x-5=6 add 5 on both sides x-5+5=6+5 x=11

OpenStudy (anonymous):

@decentNabeel X actually equals 1, but good try

OpenStudy (decentnabeel):

@redskychicken you mean x=1 ok put x=1 3x − 5 = 2x + 6 3(1)-5=2(1)+6 3-5=2+6 -2=8 so cannot be equal.. so tell me why x=11 is not correct

OpenStudy (decentnabeel):

put x=11 3x − 5 = 2x + 6 3(11)-5=2(11)+6 33-5=22+6 28=28 so that are equal on both sides so that x=11 is the correct answer are you understand @redskychicken